# HOLMBERG IX: A UNIQUE, INFANT BUT INACTIVE GALAXY as Revealed via a Multiwavelength Approach

### 03 Apr 2026 v1

@@ Abstract
In this letter, we report a novel discovery of unique characteristics for the tidal dwarf galaxy (candidate) Holmberg IX via a multiwavelength investigation. New observations are taken for deeply mapping H emission and combined with archival/published data for comprehensively probing dust, gas, and stellar populations in this galaxy. We find in Holmberg IX a dearth of dust incompatible with its rich gas and metal; globally young stellar populations with prominent FUV but deficient and marginal H emissions, distinct from other tidal dwarf galaxies ever known. By assuming normal IMF, Holmberg IX is suggested to be born Myr ago from a bursty star formation event which then rapidly ceased, with very few stars formed in the past Myr that demarcates a lower age limit for the galactic mainbody; current star formation occurs only in outskirts, bringing a conundrum about the reason for the recent quenching in such a gas-rich environment. Contradicting general expectation for tidal dwarf galaxies hosting continuous star formation, the present quiescence implies Holmberg IX currently staying in a rarely-seen transient period. Without star formation continuing, Holmberg IX is likely transforming into a dwarf spheroidal galaxy, or oppositely into a(n) (ultra-)diffuse system which will probably dissolve in the end. Instead, if Holmberg IX possesses peculiar IMF and hosts low-mass, weak-H star formation, it is able to maintain long-term survival in its current status. On whichever evolutionary pathway in reality, Holmberg IX appears as a special case updating conventional understandings of tidal dwarf galaxies and hinting potential existence of similar analogs in the universe.
@@ I INTRODUCTION
Gravitational interactions between large galaxies are likely to strip materials out of the interacting galaxies due to tidal effect and consequently generate new stellar systems with relatively small size, usually termed as ”tidal dwarf galaxies” (TDGs, Schweizer, 1978; Mirabel et al., 1992; Barnes & Hernquist, 1992). Suffering from turbulence induced by multiple gravitational fields, TDGs are not able to assemble into steady structures and hence appear irregular in morphology. Dissimilar to dwarf spheroidal/elliptical galaxies lying in a quiescent state, TDGs are expected to continuously form stars from birth to the present time, which yields strong emission lines prominently featuring in their spectra (Duc & Mirabel, 1999).
Destinies of TDGs after cessation of star formation (hereafter denoted as ”SF”) are unclear and debatable. There have been two opposite pathways proposed to chart the evolution. Without SF continuing, some of TDGs are likely to dissolve into diffuse or filamentary systems and perhaps eventually dissipate in the universe (Hibbard & Mihos, 1995; Michel-Dansac et al., 2010; Smith et al., 2013), whereas some others are suggested with a high probability to condense and transform into long-lived compact spheroidal/elliptical systems (Bournaud & Duc, 2006; Metz & Kroupa, 2007; Dabringhausen & Kroupa, 2013). In whatever evolution scenario, TDGs can hardly be preserved in present shapes for long time if stars stop forming, yet there have not been TDGs discovered with no current SF, possibly because such a period is too short to be captured.
@@
Stellar populations in Holmberg IX have been studied through observations at optical and NIR broadbands taken with Hubble Space Telescope and large ground-based telescopes (Sabbi et al., 2008; Weisz et al., 2008; Okamoto et al., 2015). By jointly analyzing color-magnitude diagrams (CMDs) and stellar density maps, these studies have found that observable stars younger than Myr show a high level of spatial concentration; older stars have a relatively uniform distribution and a slightly increasing trend towards M81 in spatial density, comparable to those in other fields at the same distance from M81. This feature suggests the old stellar component unaffiliated with Holmberg IX but associated with the outer halo of the adjacent companion M81. Therefore, stars belonging to Holmberg IX are suggested not to evolve over 250 Myr. The upper limit Myr on age is in agreement with the time for the gravitational interaction in the M81 group Myr ago (Yun, 1999), in support of the hypothesis of the tidal origin for Holmberg IX.
However, existing optical or NIR broadband photometric diagnostics are not capable of reaching young stellar populations and thereby constraining the lower age limit, which makes current SF state unclear. Excited by high-energy photons from O-type stars with lifetime of Myr, the H emission line stands in a good position to trace SF currently occurring (Kennicutt, 1998). There have been some observations of H emission in Holmberg IX taken by telescopes with small diameters, but these studies have yielded considerable deviation in estimates of SF rate (SFR) (Miller & Hodge, 1994; Kennicutt et al., 2008; Andjelic, 2011).
@@ II.1 H Emission-Line Images
Observations of Holmberg IX with FPI were carried out on December 19–21 in 2019, by use of the 6-m Big Telescope Alt-azimuth (BTA) equipped with a prime focus multimode focal reducer SCORPIO-2 (Afanasiev & Moiseev, 2011) at the Special Astrophysical Observatory of the Russian Academy of Sciences. The total exposure time was 14400 s for 40 interferograms corresponding to different gaps between FPI plates, reaching the sensitivity limit erg s-1 cm-2 arcsec-2. Data reduction was performed in compliance with Moiseev (2015), producing a datacube with a wide-field image () multiplying a high-resolution spectrum () at the position of each pixel (0″.7/pixel). The final spatial resolution was as FWHM of the PSF convolved to the largest seeing among all of the combined frames. The H FPI map was created via Voigt-profile fitting of the emission-line spectra with , and is shown in Figure 1 .
BTA Observations of Holmberg IX in the direct-imaging mode of SCORPIO-2 were performed on December 24 in 2020, with a narrowband filter centered at 6571 Å with FWHM Å. The total exposure time was 900 s with the spatial resolution (equal to the seeing in the observation night) and the sampling scale 0″.4/pixel. A continuum image was observed with a filter having a similar bandwidth centered near H and used to create a net emission-line image. Images of spectrophotometric standard stars, taken before and after the observations of the galaxy during the same night, were used to calibrate counts into absolute energetic flux. This H narrowband image, also shown in Figure 1, was applied to flux calibration of the H FPI map.
| Waveband | |||||
| (1) | (2) | (3) | (4) | (5) | (6) |
| GALEX FUV | 1516 | 268 | 14707 | 2.293 0.106 | 0.699 0.032 |
| Swift-UVOT UVW2 | 1928 | 657 | 389667 | 2.678 0.074 | 0.804 0.022 |
| GALEX NUV | 2267 | 732 | 29422 | 2.972 0.082 | 0.907 0.025 |
| Swift-UVOT UVW1 | 2600 | 693 | 158330 | 3.096 0.086 | 0.918 0.025 |
| BTA H-image | 6571 | 13 | 900 | 4.335 0.252 | |
| BTA H-FPI | 6571 | 13 | 14400 | 3.907 0.117 |
Note. — Columns: (1) Name of each filter in order of effective wavelength; (2) Effective wavelength in units of Å; (3) Bandwidth in units of Å; (4) Total exposure time in units of second; (5) Flux for Holmberg IX as a whole obtained with the aperture photometry described in Section II.4, in units of mJy for the UV data, and 10-14 erg s-1 cm-2 for the H data; (6) Flux for the area of the effective H emission in Holmberg IX obtained with the selective photometry described in Section II.4, in units of mJy for the UV data, and 10-14 erg s-1 cm-2 for the H data.
@@ II.2 UV Archival Images
Holmberg IX was imaged at UV bands with deep exposure by the Galaxy Evolution Explorer (GALEX; Lee et al., 2011) and the Swift UV/Optical Telescope (Swift-UVOT; Hoversten et al., 2011). In our study, four-band UV images were adopted, among which FUV, NUV were downloaded from Guest Investigators (GI) database of GALEX at the Multimission Archive at Space Telescope Science Institute (MAST) website,111http://galex.stsci.edu/ and UVW2, UVW1 were retrieved from the Swift-UVOT survey at the SkyView virtual observatory.222https://skyview.gsfc.nasa.gov/ Basic information on the UV and H observations including effective wavelength, bandwidth, and exposure time is listed in Table 1.
@@ II.3 Image Processing
Global background in the GALEX and Swift-UVOT images was estimated by polynomial fitting along both spatial axes in source-masked images with fields, and then subtracted from the original images (the background fitting method is addressed in Mao et al., 2014, and relevant articles therein). In order to carry out spatially resolved measurements, unification of PSFs is a requisite. In this work, the BTA, GALEX, and Swift-UVOT images were convolved to match the GALEX NUV PSF (FWHM , the lowest resolution amongst the images), with adapted kernels offered in Aniano et al. (2011). The PSF-matched images were registered on the same scale () and the same coordinate by using the SWarp software (Bertin et al., 2002).
@@
Figure 1 is an 8-stamp atlas of Holmberg IX imaged through the UV and H channels, at both of the original and convolved spatial resolutions. All measurements and analyses presented in below sections are based on the convolved images.
@@ II.4 Photometry
For the purpose of obtaining global properties of Holmberg IX, aperture photometry was utilized to measure the main body of the galaxy as a whole, by employing an elliptical aperture centered at R.A. = , Dec. = , with the diameters (corresponding to the physical scale 2.6 kpc 2.2 kpc) and the position angle . The shape of the aperture was determined manually in order to enclose the bulk of galactic brightness but avoid involving the low-brightness edge of the galaxy contaminated by background sources and foreground stars.333Foreground stars in this sky field, marked with yellow circles in Figures 1–3, are identified in optical and NIR images obtained via the Beijing-Arizona-Taipei-Connecticut multi-color wide-field survey (BATC; Fan et al., 1996; Sun et al., 2005), the Sloan Digital Sky Survey (SDSS; York et al., 2000; Alam et al., 2015), and the Two Micron All Sky Survey (2MASS; Jarrett et al., 2000, 2003). This photometric aperture is displayed as the cyan ellipse in Figure 1.
@@
In addition to the photometry on a galactic scale, subregions of the galaxy on a smaller scale were also measured in this work. The subregions were sampled in the GALEX FUV image and categorized into two classes. Emission peaks above 5 level in the FUV image were detected with the SExtractor software (Bertin & Arnouts, 1996) and defined as FUV clusters, representing the youngest stellar population inside Holmberg IX. All the other subregions than the FUV clusters were defined as diffuse regions and presumed to have evolved into older stages.
@@
Uncertainties assigned to the photometry were estimated as a quadratic sum of background deviation and calibration uncertainties in the relevant images. The background deviation is calculated as standard deviation at the non-signal area in each background-subtracted image. The calibration uncertainties taken into the quadrature are 0.05 and 0.03 mag for GALEX FUV and NUV respectively (Morrissey et al., 2007), 0.03 mag for Swift-UVOT bands (Breeveld et al., 2011), 3% for both of the SCORPIO-2 modes of the BTA observations.
@@ III.1 Dustless Trait Disclosed by IR Inspection
The sky field of Holmberg IX has been covered by a number of IR observations including the Enhanced Resolution Galaxy Atlas (WERGA) project of the Wide-field Infrared Survey Explorer (WISE; Jarrett et al., 2013), the Local Volume Legacy (LVL) survey of the Spitzer Space Telescope (Spitzer; Dale et al., 2009), and the Very Nearby Galaxies Survey (VNGS) of the Herschel Space Observatory (Herschel; Bendo et al., 2010). We visually inspect the images at WISE 22 m, Spitzer-MIPS 24 m, Herschel-PACS 70, 160 m, and Herschel-SPIRE 250, 350, 500 m bands, and find no clustering signal but only random fluctuation throughout the Holmberg IX field. Through tentative photometry in each of the IR images, we find the difference between the average values for the pixels inside and outside the elliptical aperture smaller than the background deviation. This visual plus tentatively photometric test reflects very trivial dust content (if any) in Holmberg IX with the total amount of either hot or cold dust below detection limits of the IR telescopes, coincident with previous studies of this galaxy concluding undetectable IR emission (Dale et al., 2017) and negligible dust attenuation (Hoversten et al., 2011).444Dale et al. (2017) failed to obtain integrated fluxes for Holmberg IX above detection limits at 14 IR bands from 5.8–850 m; Hoversten et al. (2011) attempted to quantify internal dust attenuation of Holmberg IX without IR data, but resulted in an estimate less than the error in measurement. On a larger scale in this sky field of the IR images, detectable IR signal is supposed to originate from foreground Galactic cirrus belonging to the Milky Way (Davies et al., 2010) or background residues unrelated with local sources. In order to corroborate the peculiarly dustless trait for Holmberg IX from a more definitive perspective, we perform a systematic analysis by referencing relevant previous studies, which is presented below.
@@
Dale et al. (2023, hereafter denoted as D23) have conducted a statistical study of spectral energy distributions (SEDs) for 258 nearby galaxies surveyed in the Local Volume Legacy program (LVL; Dalcanton et al., 2009), with Holmberg IX contained in the sample. Mid- and Far-IR images for LVL galaxies have been obtained with Spitzer-MIPS observations and measured with large apertures.555Although D23 provides IR luminosities for Holmberg IX observed with Spitzer-MIPS, the extracted photometry is likely to contain large contributions from other sources out of the galaxy, such as foreground Galactic cirrus or background sources as explained above (Nonetheless, this is merely our suspicion without solid evidence at this moment. Studies based on deeper and more resolved observations are required for drawing robust conclusions). In this case, we consider the D23 data as upper limits on the integrated IR luminosities to assess maximal dust content for Holmberg IX. In D23, dust mass for each of the galaxies has been obtained through SED fits with thoughtful modeling of dust emission, where for Holmberg IX, corresponding to the mass surface density .
@@
By quoting spectroscopic observations with the Gemini North telescope in Croxall et al. (2009), Moustakas et al. (2010) have derived the average oxygen abundance () for 2 H ii regions in Holmberg IX, from the strong-line index with the Kobulnicky & Kewley (2004, hereafter denoted as KK04) theoretical calibration on the basis of photoionization model. This metellicity corresponds to the dust-to-gas ratio according to the dust-metal relation (Galliano et al., 2018; Hunter et al., 2024). By taking the above HI-gas content into account, the expected surface density of dust mass is , much higher than the above SED-fit estimate.777Discrepancy in different metallicity diagnostics is a ubiquitous problem. From the same , the average oxygen abundance is derived alternatively as () with the McGaugh (1991) theoretical calibration but () with the Pilyugin & Thuan (2005) empirical calibration. The former is nearly the same with the KK04 estimate within the error, whereas the later is much lower and leads to the dust-to-gas ratio –, but the far smaller dust amount than the metallicity reference for Holmberg IX remains conclusive in this case.
It is intriguing that the metallicity for Holmberg IX is quite close to that for the galaxy M81 and slightly higher than that for the M81 outer disk on average. The same -KK04 prescription yields the average oxygen abundance for M81 () in Moustakas et al. (2010); for H ii regions in M81 at large radii () () in Moustakas et al. (2010) and () in Patterson et al. (2012). The coincidence of Holmberg IX with M81 in metallicity supports the tidal origin of Holmberg IX that inherits the metallicity of its parent galaxy M81 from which it was ripped apart. However, it is puzzling that Holmberg IX contains so little (or even no) dust, inconsistent with its high metallicity. Even if the IR observations were possibly not deep enough to capture effective signal, the dust content for Holmberg IX evidently lies far below the normal level for galaxies at this metallicity. The dustless albeit metal-rich property undoubtedly makes Holmberg IX an extraordinary galaxy in the local and distant universe, and deserves to explore in further depth. According to this result, we consider emissions from Holmberg IX unobscured by internal dust and thus no correction for dust attenuation is performed to the photometric fluxes in this work.

### 16 Apr 2026 v2

@@ Abstract
In this Letter, we report a novel discovery of unique characteristics for the tidal dwarf galaxy (candidate) Holmberg IX via a multiwavelength investigation. New observations are taken for deeply mapping H emission and combined with archival/published data for comprehensively probing dust, gas, and stellar populations in this galaxy. We find in Holmberg IX a dearth of dust incompatible with its rich gas and metal; globally young stellar populations with prominent far-ultraviolet but deficient and marginal H emissions, distinct from other tidal dwarf galaxies ever known. By assuming a normal initial mass function (IMF), Holmberg IX is suggested to be born Myr ago from a bursty star formation event, which then rapidly ceased, with very few stars formed in the past Myr that demarcates a lower age limit for the galactic mainbody; current star formation occurs only in outskirts, bringing a conundrum about the reason for the recent quenching in such a gas-rich environment. Contradicting the general expectation for tidal dwarf galaxies hosting continuous star formation, the present quiescence implies Holmberg IX currently staying in a rarely-seen transient period. Without star formation continuing, Holmberg IX is likely transforming into a dwarf spheroidal galaxy, or oppositely into a(n) (ultra-)diffuse system which will probably dissolve in the end. Instead, if Holmberg IX possesses peculiar IMF and hosts low-mass, weak-H star formation, it is able to maintain long-term survival in its current status. On whichever evolutionary pathway in reality, Holmberg IX appears as a special case updating conventional understandings of tidal dwarf galaxies and hinting potential existence of similar analogs in the Universe.
@@ I INTRODUCTION
Gravitational interactions between large galaxies are likely to strip materials out of the interacting galaxies due to tidal effect and consequently generate new stellar systems with relatively small size, usually termed as ”tidal dwarf galaxies” (TDGs, Schweizer, 1978; Barnes & Hernquist, 1992; Mirabel et al., 1992). Suffering from turbulence induced by multiple gravitational fields, TDGs are not able to assemble into steady structures and hence appear irregular in morphology. Dissimilar to dwarf spheroidal/elliptical galaxies lying in a quiescent state, TDGs are expected to continuously form stars from birth to the present time, which yields strong emission lines prominently featuring in their spectra (Duc & Mirabel, 1999).
Destinies of TDGs after a cessation of star formation (hereafter denoted as ”SF”) are unclear and debatable. There have been two opposite pathways proposed to chart the evolution. Without SF continuing, some of TDGs are likely to dissolve into diffuse or filamentary systems and perhaps eventually dissipate in the Universe (Hibbard & Mihos, 1995; Michel-Dansac et al., 2010; Smith et al., 2013), whereas some others are suggested with a high probability to condense and transform into long-lived compact spheroidal/elliptical systems (Bournaud & Duc, 2006; Metz & Kroupa, 2007; Dabringhausen & Kroupa, 2013). In whatever evolution scenario, TDGs can hardly be preserved in present shapes for long time if stars stop forming, yet there have not been TDGs discovered with no current SF, possibly because such a period is too short to be captured.
@@
Stellar populations in Holmberg IX have been studied through observations at optical and near-infrared (NIR) broadbands taken with Hubble Space Telescope and large ground-based telescopes (Sabbi et al., 2008; Weisz et al., 2008; Okamoto et al., 2015). By jointly analyzing color-magnitude diagrams (CMDs) and stellar density maps, these studies have found that observable stars younger than Myr show a high level of spatial concentration; older stars have a relatively uniform distribution and a slightly increasing trend toward M81 in spatial density, comparable to those in other fields at the same distance from M81. This feature suggests the old stellar component unaffiliated with Holmberg IX but associated with the outer halo of the adjacent companion M81. Therefore, stars belonging to Holmberg IX are suggested not to evolve over 250 Myr. The upper limit Myr on age is in agreement with the time for the gravitational interaction in the M81 group – Myr ago (Yun, 1999), in support of the hypothesis of the tidal origin for Holmberg IX.
However, existing optical or NIR broadband photometric diagnostics are not capable of reaching young stellar populations and thereby constraining the lower age limit, which makes current SF state unclear. Excited by high-energy photons from O-type stars with lifetimes of Myr, the H emission line stands in a good position to trace SF currently occurring (Kennicutt, 1998, hereafter denoted as K98). There have been some observations of H emission in Holmberg IX taken by telescopes with small diameters, but these studies have yielded considerable deviation in estimates of SF rate (SFR; Miller & Hodge, 1994; Kennicutt et al., 2008; Andjelic, 2011).
@@
In the following content, we describe the H observations and compilation of the ancillary data in Section II, present the results of data analysis in Section III, and discuss the implication of the results in Section V.
| Waveband | |||||
| (1) | (2) | (3) | (4) | (5) | (6) |
| GALEX FUV | 1516 | 268 | 14,707 | 2.293 0.106 | 0.699 0.032 |
| Swift-UVOT UVW2 | 1928 | 657 | 389,667 | 2.678 0.074 | 0.804 0.022 |
| GALEX NUV | 2267 | 732 | 29,422 | 2.972 0.082 | 0.907 0.025 |
| Swift-UVOT UVW1 | 2600 | 693 | 158,330 | 3.096 0.086 | 0.918 0.025 |
| BTA H-image | 6571 | 13 | 900 | 4.335 0.252 | |
| BTA H-FPI | 6571 | 13 | 14,400 | 3.907 0.117 |
Note. — Columns: (1) Name of each filter in order of effective wavelength; (2) Effective wavelength in units of Å; (3) Bandwidth in units of Å; (4) Total exposure time in units of seconds; (5) Flux for Holmberg IX as a whole obtained with the aperture photometry described in Section II.4, in units of mJy for the UV data, and 10-14 erg s-1 cm-2 for the H data; (6) Flux for the area of the effective H emission in Holmberg IX obtained with the selective photometry described in Section II.4, in units of mJy for the UV data, and 10-14 erg s-1 cm-2 for the H data.
@@ II.1 H Emission-line Images
Observations of Holmberg IX with FPI were carried out on December 19–21 in 2019, by use of the 6 m Big Telescope Alt-azimuth (BTA) equipped with a prime focus multimode focal reducer SCORPIO-2 (Afanasiev & Moiseev, 2011) at the Special Astrophysical Observatory of the Russian Academy of Sciences. The total exposure time was 14,400 s for 40 interferograms corresponding to different gaps between FPI plates, reaching the sensitivity limit erg s-1 cm-2 arcsec-2. Data reduction was performed in compliance with Moiseev (2015), producing a datacube with a wide-field image () multiplying a high-resolution spectrum () at the position of each pixel (0″.7 pixel-1). The final spatial resolution was as FWHM of the point-spread function (PSF) convolved to the largest seeing among all of the combined frames. The H FPI map was created via Voigt-profile fitting of the emission-line spectra with the signal-to-noise ratio , and is shown in Figure 1 .
BTA Observations of Holmberg IX in the direct-imaging mode of SCORPIO-2 were performed on December 24 in 2020, with a narrowband filter centered at 6571 Å with FWHM Å. The total exposure time was 900 s with the spatial resolution (equal to the seeing in the observation night) and the sampling scale 0″.4 pixel-1. A continuum image was observed with a filter having a similar bandwidth centered near H and used to create a net emission-line image. Images of spectrophotometric standard stars, taken before and after the observations of the galaxy during the same night, were used to calibrate counts into absolute energetic flux. This H narrowband image, also shown in Figure 1, was applied to flux calibration of the H FPI map.
@@ II.2 UV Archival Images
Holmberg IX was imaged at ultraviolet (UV) bands with deep exposure by the Galaxy Evolution Explorer (GALEX; Lee et al., 2011) and the Swift UV/Optical Telescope (Swift-UVOT; Hoversten et al., 2011). In our study, four-band UV images were adopted, among which far-UV (FUV), near-UV (NUV) were downloaded from Guest Investigators database of GALEX at the Multimission Archive at Space Telescope Science Institute (MAST) website,111http://galex.stsci.edu/ and UVW2, UVW1 were retrieved from the Swift-UVOT survey at the SkyView virtual observatory.222https://skyview.gsfc.nasa.gov/ Basic information on the UV and H observations including effective wavelength, bandwidth, and exposure time is listed in Table 1.
@@ II.3 Image Processing
Global background in the GALEX and Swift-UVOT images was estimated by polynomial fitting along both spatial axes in source-masked images with fields, and then subtracted from the original images (the background fitting method is addressed in Mao et al., 2014, and relevant articles therein). In order to carry out spatially resolved measurements, unification of PSFs is a requisite. In this work, the BTA, GALEX, and Swift-UVOT images were convolved to match the GALEX NUV PSF (FWHM , the lowest resolution amongst the images), with adapted kernels offered in Aniano et al. (2011). The PSF-matched images were registered on the same scale (1″.5 pixel-1) and the same coordinate by using the SWarp software (Bertin et al., 2002).
@@
Figure 1 is an eight-stamp atlas of Holmberg IX imaged through the UV and H channels, at both of the original and convolved spatial resolutions. All measurements and analyses presented in below sections are based on the convolved images.
@@ II.4 Photometry
For the purpose of obtaining global properties of Holmberg IX, aperture photometry was utilized to measure the main body of the galaxy as a whole, by employing an elliptical aperture centered at R.A. = , Decl. = , with the diameters (corresponding to the physical scale 2.6 kpc 2.2 kpc) and the position angle . The shape of the aperture was determined manually in order to enclose the bulk of galactic brightness but avoid involving the low-brightness edge of the galaxy contaminated by background sources and foreground stars.333Foreground stars in this sky field, marked with yellow circles in Figures 1–3, are identified in optical and NIR images obtained via the Beijing-Arizona-Taipei-Connecticut multicolor wide-field survey (i.e., BATC; Fan et al., 1996; Sun et al., 2005), the Sloan Digital Sky Survey (i.e., SDSS; York et al., 2000; Alam et al., 2015), and the Two Micron All Sky Survey (i.e., 2MASS; Jarrett et al., 2000, 2003). This photometric aperture is displayed as the cyan ellipse in Figures 11, the top panels of Figure 2, and the left panel of Figure 3.
@@
In addition to the photometry on a galactic scale, subregions of the galaxy on a smaller scale were also measured in this work. The subregions were sampled in the GALEX FUV image and categorized into two classes. Emission peaks above level in the FUV image were detected with the SExtractor software (Bertin & Arnouts, 1996) and defined as FUV clusters, representing the youngest stellar population inside Holmberg IX. All the other subregions than the FUV clusters were defined as diffuse regions and presumed to have evolved into older stages.
@@
Uncertainties assigned to the photometry were estimated as a quadratic sum of background deviation and calibration uncertainties in the relevant images. The background deviation is calculated as standard deviation at the non-signal area in each background-subtracted image. The calibration uncertainties taken into the quadrature are 0.05 and 0.03 mag for GALEX FUV and NUV, respectively (Morrissey et al., 2007), 0.03 mag for Swift-UVOT bands (Breeveld et al., 2011), 3% for both of the SCORPIO-2 modes of the BTA observations.
@@ III.1 Dustless Trait Disclosed by IR Inspection
The sky field of Holmberg IX has been covered by a number of IR observations including the Enhanced Resolution Galaxy Atlas project of the Wide-field Infrared Survey Explorer (WISE; Jarrett et al., 2013), the Local Volume Legacy (LVL) survey of the Spitzer Space Telescope (Spitzer; Dale et al., 2009), and the Very Nearby Galaxies Survey of the Herschel Space Observatory (Herschel; Bendo et al., 2010). We visually inspect the images at WISE 22 m, Spitzer-MIPS 24 m, Herschel-PACS 70, 160 m, and Herschel-SPIRE 250, 350, 500 m bands, and find no clustering signal but only random fluctuation throughout the Holmberg IX field. Through tentative photometry in each of the IR images, we find the difference between the average values for the pixels inside and outside the elliptical aperture smaller than the background deviation. This visual plus tentatively photometric test reflects very trivial dust content (if any) in Holmberg IX with the total amount of either hot or cold dust below detection limits of the IR telescopes, coincident with previous studies of this galaxy concluding undetectable IR emission (Dale et al., 2017) and negligible dust attenuation (Hoversten et al., 2011).444Dale et al. (2017) failed to obtain integrated fluxes for Holmberg IX above detection limits at 14 IR bands from 5.8 to 850 m; Hoversten et al. (2011) attempted to quantify internal dust attenuation of Holmberg IX without IR data, but resulted in an estimate less than the error in measurement. On a larger scale in this sky field of the IR images, detectable IR signals are supposed to originate from foreground Galactic cirrus belonging to the Milky Way (Davies et al., 2010) or background residues unrelated with local sources. In order to corroborate the peculiarly dustless trait for Holmberg IX from a more definitive perspective, we perform a systematic analysis by referencing relevant previous studies, which is presented below.
@@
Dale et al. (2023, hereafter denoted as D23) have conducted a statistical study of spectral energy distributions (SEDs) for 258 nearby galaxies surveyed in the LVL program (Dalcanton et al., 2009), with Holmberg IX contained in the sample. Mid- and far-IR images for LVL galaxies have been obtained with Spitzer-MIPS observations and measured with large apertures.555Although D23 provides IR luminosities for Holmberg IX observed with Spitzer-MIPS, the extracted photometry is likely to contain large contributions from other sources out of the galaxy, such as foreground Galactic cirrus or background sources as explained above (Nonetheless, this is merely our suspicion without solid evidence at this moment; studies based on deeper and more resolved observations are required for drawing robust conclusions). In this case, we consider the D23 data as upper limits on the integrated IR luminosities to assess maximal dust content for Holmberg IX. In D23, dust mass for each of the galaxies has been obtained through SED fits with thoughtful modeling of dust emission, where for Holmberg IX, corresponding to the mass surface density .
@@
By quoting spectroscopic observations with the Gemini North telescope in Croxall et al. (2009), Moustakas et al. (2010) have derived the average oxygen abundance () for two H ii regions in Holmberg IX, from the strong-line index with the Kobulnicky & Kewley (2004, hereafter denoted as KK04) theoretical calibration on the basis of photoionization model. This metellicity corresponds to the dust-to-gas ratio according to the dust-metal relation (Galliano et al., 2018; Hunter et al., 2024). By taking the above H i-gas content into account, the expected surface density of dust mass is , much higher than the above SED-fit estimate.777Discrepancy in different metallicity diagnostics is a ubiquitous problem. From the same , the average oxygen abundance is derived alternatively as () with the McGaugh (1991) theoretical calibration but () with the Pilyugin & Thuan (2005) empirical calibration. The former is nearly the same with the KK04 estimate within the error, whereas the latter is much lower and leads to the dust-to-gas ratio –, but the far smaller dust amount than the metallicity reference for Holmberg IX remains conclusive in this case.
It is intriguing that the metallicity for Holmberg IX is quite close to that for the galaxy M81 and slightly higher than that for the M81 outer disk on average. The same -KK04 prescription yields the average oxygen abundance for M81 () in Moustakas et al. (2010); for H ii regions in M81 at large radii () () in Moustakas et al. (2010) and () in Patterson et al. (2012). The coincidence of Holmberg IX with M81 in metallicity supports the tidal origin of Holmberg IX that inherits the metallicity of its parent galaxy M81 from which it was ripped apart. However, it is puzzling that Holmberg IX contains so little (or even no) dust, inconsistent with its high metallicity. Even if the IR observations were possibly not deep enough to capture effective signals, the dust content for Holmberg IX evidently lies far below the normal level for galaxies at this metallicity. The dustless albeit metal-rich property undoubtedly makes Holmberg IX an extraordinary galaxy in the local and distant Universe, and deserves to be explored in further depth. According to this result, we consider emissions from Holmberg IX unobscured by internal dust, and thus, no correction for dust attenuation is performed to the photometric fluxes in this work.